OpenStudy (erniellyd28):

(to evaluate a function) given f(x)=3x-4,find f(4)

OpenStudy (anonymous):

plug in 4 for x

OpenStudy (anonymous):

3(4)-4=12-4=8

OpenStudy (erniellyd28):

thats all there is to it?

OpenStudy (anonymous):

yep!

OpenStudy (erniellyd28):

damn im a dork lol. thank u so much for your help .math has always been my weakness

OpenStudy (anonymous):

its ok! :)

OpenStudy (erniellyd28):

if its ok and if your not busy can i ask your help on one more problem please?

OpenStudy (anonymous):

well let me look at it first, then i will If i can help or not!

OpenStudy (anonymous):

..............then i will tell if i.......

OpenStudy (erniellyd28):

(find the x- and y- intercepts) 3x+4y=12

OpenStudy (anonymous):

Plug in 0 for x to find y-int plug in 0 for y to find x-int

OpenStudy (erniellyd28):

i dont have a calculator at all much less one with those functions

OpenStudy (erniellyd28):

thank u so much for all your help anyway i :) i will figure this one out on my own

OpenStudy (anonymous):

x-int: 3x+4(0) = 12 3x=12 x=4-------------->c-int y-int 3(0)+4y=12 4y=12 y=3

OpenStudy (anonymous):

(4,0) (0,3)

OpenStudy (erniellyd28):

again thank u so very much so sorry to have made u go through all that trouble :) u have a good night

OpenStudy (anonymous):

its ok! take care! nite!!! :)